OpenStudy (anonymous):

An airplane leaves an airport and flies due west 150 miles and then 170 miles in the direction S 49.17°W. How far is the plane from the airport (round to the nearest mile)?

OpenStudy (anonymous):

Please help me

OpenStudy (anonymous):

someone

OpenStudy (anonymous):

anyone

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (anonymous):

Let x= distance from airport to plane in miles x^2=170^2+150^2-2(170)(150)cos(49.17+90) x^2=89,989.29 x=300 miles

OpenStudy (anonymous):

i dont think its right

OpenStudy (anonymous):

@goformit100 are you still there?